how can we identify no2 and br2 seperately
1
Expert's answer
2012-09-14T09:49:37-0400
Try to dissolve it in the water. In the case of NO2 during dissolving it will
loose color (from red to uncolored)
NO2 + H2O= HNO3 + HNO2
In the case of
Br2 after dissolving in the first period of time it will be red solution,but
after some time it will be uncolored too
Br2+H2O=HBr + HBrO